Solve the equation 8sin^2𝑥−7=2cos𝑥 for 0≤𝑥≤360°, giving your answers to 1 decimal place.


1
Expert's answer
2022-03-17T18:45:09-0400

Explanations & Calculations


"\qquad\qquad
\begin{aligned}
\small 8\sin^2x-7&=\small 2\cos x\\
\small 8(1-\cos^2x)-7&=\small 2\cos x\\
\small 8\cos^2x+2\cos x-1&=\small 0\\
\small (4\cos x-1)(2\cos x+1)&=\small 0\\
\small \cos x &= \small \frac{1}{4}\quad: \, x = 75.5^0\,\&\,284.5^0\\
\small or\\
\small \cos x&=\small -\frac{1}{2}\quad: x=120.0^0\,\&\,240.0^0
\end{aligned}"